What is the simple interest on a Php 500 loan at 25% interest to be paid at the end of 6 months

1
Expert's answer
2021-12-14T12:24:06-0500

"Simple\space interest

= Amount \space of\space loan \times Interest\space rate \% \times time\\


= 500 \times 25\% \times \frac{6}{12}\\

=62.50"